1/2x-5/3=-1/2x+19/4
Answer by checkley77(12844)

1/2x-5/3=-1/2x+19/4
X/2+X/2=19/4+5/3
X=(19*3+5*4)/(4*3)
X=(57+20)/12
X=77/12 ANS.
PROOF:
1/2*77/12-5/3=-1/2*77/12+19/4
77/24-5/3=-77/24+19/4
77/24+77/24=19/4+5/3
154/24=(19*3+5*4)/12
154/24=(57+20)/12
154/24=77/12
24*77=12*154
1848=1848